OpenStudy (anonymous):

Why can't sqrt(x^4 - 4x^2) = x^2 - 4x?

jimthompson5910 (jim_thompson5910):

because sqrt(a+b) does not equal sqrt(a)+sqrt(b)

OpenStudy (anonymous):

Just try squaring x² - 4x (x² - 4x)² = x^4 - 8x³ + 16x² ≠ x^4 - 4x²

jimthompson5910 (jim_thompson5910):

and if it did, then sqrt(4x^2) = 2x and not 4x